# Beyond Prompting: An Efficient Embedding Framework for Open-Domain Question Answering

### v2

@@ 3.2 Retriever: Self-Refinement Driven Reranking
To learn $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, we adopt an unsupervised contrastive loss Oord et al. (2018) that encourages $\mathbf{e}_{q_{\text{new}}}$ to focus on passages containing the candidate answers from ${y}$. Specifically, we treat passages that contain at least one candidate as positive, and those that do not as negative. Because negatives are significantly more abundant, we sample them at a fixed ratio of 5:1 relative to positives in each training batch to maintain balance. Once the parameters are updated, we use $\mathbf{e}_{q_{\text{new}}}$ to re-query the retriever, effectively re-ranking the knowledge base $\mathcal{B}$ so that passages with correct answers appear more prominently. This addresses the inefficiency of multi-turn prompt-level reranking and increases the likelihood that the subsequent Reader module receives high-quality evidence.
@@ 4.1 Setups
Model Retriever & Framework Dataset
@@ 4.2 Main Results
Dataset Method Time/query (min) ↓ Tokens /query ↓
@@ 5 Analysis
Retriever& Rerank Framework HotpotQA 2Wiki NQ WebQ
Avg. GT @Top-10 Time /Query(s) Avg. GT @Top-10 Time /Query(s) Avg. GT @Top-10 Time /Query(s) Avg. GT @Top-10 Time /Query(s)
@@
Retriever & Rerank Module HotpotQA 2Wiki NQ WebQ
@@ A.1 Answer Candidates Generation
Each answer should be in the form: (a) xx, (b) yy, and should not exceed 3 words."
@@ Appendix B Theoretical Discussion on Embedding Space
Considering the space between two adjacent variables in order statistics $\Delta_{i}=u_{(i)}-u_{(i+1)}$ with the only large half where $21\leq i\leq k/$ 2, we have:
@@
[…]
@@ Appendix D Further Fusion Study on Contrastive Learning and the Embedding Refinement
New Experiment EmbQA w/o exploratory & contrastive learning in rerank 33.8 / 50.5 22.4 / 31.4 40.2 / 53.0 34.0 / 49.1 32.6 / 46.0

### v3

@@ 3.2 Retriever: Self-Refinement Driven Reranking
To learn $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, we adopt an unsupervised contrastive loss Oord et al. (2018) that encourages $\mathbf{e}_{q_{\text{new}}}$ to focus on passages containing the candidate answers from y{y}. Specifically, we treat passages that contain at least one candidate as positive, and those that do not as negative. Because negatives are significantly more abundant, we sample them at a fixed ratio of 5:1 relative to positives in each training batch to maintain balance. Once the parameters are updated, we use $\mathbf{e}_{q_{\text{new}}}$ to re-query the retriever, effectively re-ranking the knowledge base $\mathcal{B}$ so that passages with correct answers appear more prominently. This addresses the inefficiency of multi-turn prompt-level reranking and increases the likelihood that the subsequent Reader module receives high-quality evidence.
@@ 4.1 Setups
Model
Retriever &
Framework
Dataset
@@ 4.2 Main Results
Dataset Method
Time/query (min) ↓
Tokens /query ↓
@@ 5 Analysis
Retriever&
Rerank Framework
HotpotQA 2Wiki NQ WebQ
Avg. GT
@Top-10
Time
/Query(s)
Avg. GT
@Top-10
Time
/Query(s)
Avg. GT
@Top-10
Time
/Query(s)
Avg. GT
@Top-10
Time
/Query(s)
@@
Retriever
& Rerank Module
HotpotQA 2Wiki NQ WebQ
@@ A.1 Answer Candidates Generation
Each answer should be in the form: (a) $x$, (b) $y$, and should not exceed 3 words."
@@ Appendix B Theoretical Discussion on Embedding Space
Considering the space between two adjacent variables in order statistics $\Delta_{i}=u_{(i)}-u_{(i+1)}$ with the only large half where $1\leq i\leq k/2$, we have:
@@
$\mathbb{E}(\Delta_{i}^{2})\leq\frac{2}{i^{2}}\mathbb{E}(e^{\frac{u_{(i)}^{2}}{2}}\int_{u_{(i)}}^{+\infty}e^{-\frac{t^{2}}{2}}dt).$
Obviously, there are three parts in the expectation of space $\Delta_{i}$ in the order statistics $u_{(i)}$. With the increasing of order index $i$ in $u_{(i)}$, the value of $u_{(i)}$ is decreasing. Thus, the $2/i^{2}$ will converge to zero, and the experiential term $e^{u_{i}^{2}/2}$ will decrease to 1 since most embeddings in LLM are around the original point in current popular models like LLama or Qwen. For the third term of $\int_{u_{(}i)}^{+\infty}e^{-\frac{t^{2}}{2}}dt$, it is a survival function from a Gaussian distribution, then it should be bounded by 1. (In fact, it should be bounded by 0.5 if we only consider the positive $u_{(i)}$ from the first half of the order statistics)
@@ Appendix D Further Fusion Study on Contrastive Learning and the Embedding Refinement
New Experiment
EmbQA w/o exploratory &
contrastive learning in rerank
33.8 / 50.5 22.4 / 31.4 40.2 / 53.0 34.0 / 49.1 32.6 / 46.0